# Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations

### v2

@@ 7 Conclusion
Whereas most existing interpretability methods highlight important tokens in the model’s output [63, 64], gSMILE shifts attention to the input space. By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, LLaMA, Claude.ai, and Gemma. The resulting heatmaps and attribution scores provide an intuitive understanding of model behavior, empowering users to control and refine prompt design better. gSMILE was systematically evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including fidelity, stability, and consistency.

### v4

@@ 7 Conclusion
Whereas most existing interpretability methods highlight essential tokens in the model’s output [64, 65], gSMILE shifts attention to the input space. By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, Meta’s LLaMA 3.1 Instruct Turbo (70B), Claude.ai, and Gemma. The resulting heatmaps and attribution scores provide an intuitive understanding of model behaviour, empowering users to control and refine prompt design better. gSMILE was systematically evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including fidelity, stability, and consistency.